Let x₁,x₂,..., Xn be a random sample from N(μ,sigma²), where u is known. Construct a (1-a) * 100% Confidence Interval for sigma².


1
Expert's answer
2021-10-29T04:38:59-0400

For building confidence interval we will determine next random value:

"H = {\frac{(x{\scriptscriptstyle1}-u)^{2}+...+(x{\scriptscriptstyle n}-u)^{2}} {σ^{2}}}" ~ "Xi^{2}(n)", where u is μ(mean value)

Let "{(x{\scriptscriptstyle1}-u)^{2}+...+(x{\scriptscriptstyle n}-u)^{2}} = sum"

Since we have 1 - a confidence interval, then we have to find such m and n, that

"P(m<H<n)=1-a"

Technically, we can pick up any "m = Xi^{2}(k{\scriptscriptstyle 1},n), n = Xi^{2}(1-k{\scriptscriptstyle 2},n)" , where "k{\scriptscriptstyle 1}+k{\scriptscriptstyle 2}=a"

For example, we can put "k{\scriptscriptstyle 1}={\frac a 2}, k{\scriptscriptstyle 2}=1-{\frac a 2}" . Now we have:

"P(Xi^{2}({\frac a 2},n)<H<Xi^{2}(1-{\frac a 2},n))=1-a"

"P(Xi^{2}({\frac a 2},n)<{\frac{sum} {σ^{2}}} <Xi^{2}(1-{\frac a 2},n))=1-a"

"P({\frac {sum} {Xi^{2}({\frac a 2},n)}}<σ^{2} <{\frac {sum} {Xi^{2}(1-{\frac a 2},n)}})=1-a"

The 1-a confidence interval is for "\u03c3^{2}" is "({\frac {sum} {Xi^{2}({\frac a 2},n)}},{\frac {sum} {Xi^{2}(1-{\frac a 2},n)}})"